OpenStudy (anonymous):

speed skaters spend many hours training on and off the ice to improve their strength and conditioning. Suppose a team has a monthly training budget of $7000. ice rental costs $125/h, and gym rental for strength training costs $55/h. Determine the solution region and all possible combination of training time that the team can afford. THANKSSS.

OpenStudy (pfenn1):

Let x = the hours of ice time and y = the hours of gym rental. Then \[125x+55y \le 7000\]

OpenStudy (anonymous):

how do you find the combinations?

OpenStudy (pfenn1):

Good question. I don't have any bright ideas. I would just set y = 0 and calculate the maximum x. and then set x= 0 and calculate the maximum y.\[125x \le7000\]\[0\le x \le56\]\[55y \le 7000\]\[0 \le y \le 127\]On a graph it would something like